OpenStudy (anonymous):

some one please teach me how to do these please? im so confused :c simplify the expression x^2 + 6x + 9 / 5x + 15

OpenStudy (jim766):

x^2 + 6x + 9 do you know how to factor this?

OpenStudy (jim766):

factoring means rewriting so it is 2 things mulitplied together ( ) ( ) what times what = x^2

OpenStudy (anonymous):

i have no clue...

OpenStudy (jim766):

(x + 3 )(x + 3) = x^2 + 6x + 9 ------------ ------------ 5x + 15 do you know the commond denominator

OpenStudy (jim766):

what divides evenly into both 5 and 15

OpenStudy (anonymous):

3

OpenStudy (anonymous):

wait how did you get that equation?!?!

OpenStudy (jim766):

which one?

OpenStudy (anonymous):

(x + 3 )(x + 3) = x^2 + 6x + 9 ------------ ------------ 5x + 15

OpenStudy (jim766):

x^2 + 6x + 9 if we factor it, it will be (x + 3)(x + 3) try it, do you know how to multiply (x+ 3)(x + 3)

OpenStudy (anonymous):

um 3x and 3x? im sorry im so confused

OpenStudy (jim766):

ok lets start from the begining. (x + 3)(x+ 3) have your heard of FOIL?

OpenStudy (anonymous):

ive heard of it.. but i dont how to use it

OpenStudy (jim766):

It tells us how to multiply to binomials (x + 3) (x + 3 ) F irst O uter I nner L asts x *x x * 3 3 * x 3 * 3 do you see where I got this?

OpenStudy (anonymous):

Ohhh i see! but i still dont get who you got the 3.. do you use 3 in every expression like this one?

OpenStudy (anonymous):

how not who **

OpenStudy (jim766):

ok just lets finish this F irst O uter I nner L asts x *x x * 3 3 * x 3 * 3 x^2 + 3x + 3x + 9 x^2 + 6x + 9

OpenStudy (jim766):

x^2 + 6x + 9 now ,do you see how this is in your original problem?

OpenStudy (anonymous):

OHHHH yes! thank you! okay go on :)

OpenStudy (anonymous):

oh wait you factor the numberator

OpenStudy (anonymous):

okay i got that part and then you factor the denominator which is....?

OpenStudy (jim766):

x^2 + 6x + 9 to factor this, I start off with what time what = x^2 x * x = x^2 , so I put it in the front (x )(x )

OpenStudy (jim766):

x time x = x^2

OpenStudy (anonymous):

x * 3 and x* 3 ??

OpenStudy (jim766):

(x + 3)(x +3)

OpenStudy (anonymous):

right right. sorry... go on! :)

OpenStudy (jim766):

now look at the denominator 5x + 15 what number divides evenly into both 5 and 15 (it is not 3)

OpenStudy (anonymous):

5 !

OpenStudy (jim766):

right, that is the common denominator....

OpenStudy (anonymous):

so the factors are ( x + 5) and ( x +5 ) ??

OpenStudy (anonymous):

or just ( x + 3 ) ( x + 3 ) --------------- 5

OpenStudy (jim766):

this one is a little different... 5x + 15 common denominator is 5 5( ) now divide everthing by 5

OpenStudy (jim766):

5x + 15 common denominator is 5 --- --- 5 5 5( x + 3) now divide everthing by 5

OpenStudy (jim766):

(x + 3 )(x + 3) = x^2 + 6x + 9 ------------ ------------ 5 (x + 3) 5x + 15

OpenStudy (jim766):

what can You cancel out?

OpenStudy (anonymous):

the threes !

OpenStudy (jim766):

the (x +3)s

OpenStudy (anonymous):

yes

OpenStudy (jim766):

leaving you with x + 3 ------ 5

OpenStudy (anonymous):

thats the answer??

OpenStudy (jim766):

you only can cancel one out so the other is left

OpenStudy (jim766):

yet that is the answer

OpenStudy (jim766):

* yep

OpenStudy (anonymous):

Ohhh i get it thank you sooooooo much! heres your medal :D

OpenStudy (jim766):

yw

OpenStudy (anonymous):

can you help me with another one? i just need help going through it...

OpenStudy (jim766):

sure...

OpenStudy (anonymous):

okay its \[2x^2 - 4a +2 / 3a^2 - 3\]

OpenStudy (anonymous):

so i factor the numberator which is... (x-2) (x+2) ??

OpenStudy (jim766):

yes you have to factor the numerator...but start by looking for a GCF What divides evenly into 2, -4 and 2

OpenStudy (anonymous):

its -2 or 2

OpenStudy (jim766):

yes, usually we will keep it positive, so lets use 2 2a^2 -4a + 2 --- --- --- 2 2 2 2 (

OpenStudy (jim766):

what is 2a^2 divided by 2

OpenStudy (anonymous):

a^2 ??

OpenStudy (jim766):

\[\frac{ 2a^2 }{ 2 }\]

OpenStudy (anonymous):

a squared

OpenStudy (jim766):

right, divide each of the terms by 2

OpenStudy (jim766):

2a^2 -4a + 2 2 (a^2 - 2a + 1) do you see where I got that?

OpenStudy (anonymous):

so its a^2 - 2a +1

OpenStudy (anonymous):

yes you divided 2 to every thing

OpenStudy (anonymous):

by everything i mean the numberator numbers

OpenStudy (jim766):

ok now we have to factor that 2 (a^2 -2a + 1 ) 2 (a )(a ) what 2 numbers mult to get 1 and add to get -2?

OpenStudy (anonymous):

-1 and -1

OpenStudy (jim766):

right that is what you put in the parentheses 2 (a - 1 )(a - 1 ) now that is factored

OpenStudy (anonymous):

ok

OpenStudy (jim766):

now factor 3a^2 -3 what is the GCF?

OpenStudy (jim766):

what goes evenly into 3 and -3?

OpenStudy (anonymous):

1?

OpenStudy (jim766):

yes 1 does, but is it the biggest that goes evenly into both?

OpenStudy (anonymous):

3 !!!! duhh!! haha oops!

OpenStudy (jim766):

right now divide everything by 3 3a^2 -3 -- -- 3 3 3 (

OpenStudy (anonymous):

all thats left is a^2

OpenStudy (anonymous):

a^2 -1

OpenStudy (jim766):

right... 3 (a^2 - 1) this can still be factored though...

OpenStudy (jim766):

3 (a )( a )

OpenStudy (jim766):

what 2 numbers can you mult together to get -1

OpenStudy (anonymous):

3 (a-1) ( a-1) ??

OpenStudy (jim766):

almost... -1* -1 = 1

OpenStudy (anonymous):

ohh ok i get i

OpenStudy (anonymous):

so its 3 (a +1) (a + 1)

OpenStudy (jim766):

almost again -1 * +1 = -1 2 (a - 1 )(a - 1 ) ---------------------- which cancels out? 3 (a +1) (a - 1)

OpenStudy (anonymous):

the ones cancel out so its just 2/3

OpenStudy (anonymous):

right?

OpenStudy (jim766):

anything divided by itself equal 1 (a-1)/(a-1) = 1 one (a -1) on top and one (a-1) on bottom cancels, whats left?

OpenStudy (anonymous):

2(a-1) ------ 3 (a+1)

OpenStudy (jim766):

thats right, thats the answer...

OpenStudy (anonymous):

THANKYOU so much for your help! im not going to fail! :D haha

OpenStudy (jim766):

keep working at it, youll get it!

OpenStudy (anonymous):

thanks!